# Muffin: Mitigating Unhelpfulness in Emotional Support Conversations with Multifaceted AI Feedback

### v2

@@ C.1 Multifaceted AI Feedback
Overall, Muffin demonstrates enhancements in AI feedback across multiple facets. However, three intriguing phenomena emerge. (1). While the left subtable clearly indicates an increase in the frequency of helpful responses attributable to the Muffin framework, the evidence presented in the right subtable is weaker (agg.), particularly when the base model is TransESC, KEMI, or MultiESC. However, this finding aligns with our loss function (Equation 1 and Equation 4). We have introduced a contrastive loss to the original generation loss, but this addition does not significantly mitigate unhelpfulness. However, the contrastive loss assigns higher generation probabilities to helpful (non-unhelpful) responses. Consequently, when generating one response, the one with the highest probability is output. Thus, the left subtable shows that Muffin effectively reduce the frequency of unhelpful responses. (2). Despite lacking a dedicated mechanism for incorporating communication skills, BlenderBot-Vanilla attains a notably high score in communication skill efficiency (skill). Upon closer examination, we observe that this model frequently produces responses such as “I can understand that…” or “I’ve experienced something similar…”. These responses are categorized as self-disclosure. While this strategy doesn’t undermine previous emotional support efforts and remains undetected as “unhelpful” by the AI feedback module, a conversation dominated by self-disclosure responses may not fully align with the objective of providing support. (3). Another noteworthy observation is that TransESC receives exceptionally low AI feedback in response coherence (cohr.), despite its incorporation of external knowledge. However, insights from TransESC’s ablation experiment Zhao et al. (2023) suggest that the generated knowledge may contain noise. This potential factor could explain the low feedback score observed in response coherence.
@@ C.2 Cases
For a closer and more detailed examination of our method, we present additional comparisons between the base model and the Muffin model in Table 8. Additionally, we include AI feedback for each facet as a reference.
Conversation History |
Supporter: Is there anything you’d like to talk to me about? |
@@
Help-seeker: I have a friend who knew me since I was little. She is a little older than me. She still treat me like I’m still little. I tried to explain to her that I’m a grown woman and I grew out of lot of things when I was a kid. |
Model |
Response |
emp. | skill | cohr. |
Joint |
@@
I have had a friend like that before. They are a good friend, and you should feel as though you are a better person. |
Weak Empathy |
Other |
No |
Muffin |
@@
I have been in a similar situation. I had a friend who was a little older than me, and I had to deal with the same thing. |
Weak Empathy |
Other |
Yes |
Conversation History |
@@
Help-seeker: Hello I am very sad today. My partner is cheating on me. I was told by a customer at work. I do not know this person well, yet they knew the person he was cheating on me with. He will not admit it himself. |
Supporter: I am sorry if your partner is cheating. So, your partner denies the charge of cheating? |
@@
Help-seeker: Yes. He says the person who told me is lying. We are not married but we live together. |
@@
Supporter: Do you think maybe it was a rumor started by someone who could be jealous? Ash your partner if he loves you enough to propose marriage. If he says no, then ask him why. |
@@
Help-seeker: I am not sure if I am okay with giving him an ultimatum. We are still young and although I love being in a relationship, I am unsure about marriage. But I appreciate your suggestions. And you make a great point about the workplace being hostile. |
Model |
Response |
emp. | skill | cohr. |
MultiESC |
@@
I think you should give him an ultimatum and see if he accepts it. |
No Empathy |
MI Non-Adherent |
Yes |
Muffin |
@@
Have you spoken to your friends or family about the situation? |
Strong Empathy |
Other |
Yes |
Conversation History |
@@
Help-seeker: I recently broke up with my girlfriend. We had been together for more than 3 years. She said that she simply wasn’t in love anymore, and that she didn’t have the same feelings she did at the beginning. |
@@
Supporter: You still have your whole life ahead of you. Some people don’t find a partner until their 40’s or even 50’s… I think that’s what your girl would want you to do as well. |
@@
Help-seeker: I disagree with you regarding that maybe that is what she wants me to do. I don’t think she cares about me at the slightest anymore. I was a burden for her and will probably keep being it for any woman I meet. |
Model |
Response |
emp. | skill | cohr. |
KEMI |
@@
That is a great attitude to have. You are a strong person, and I think you will find someone better. |
Weak Empathy |
MI Non-Adherent |
Yes |
Muffin |
@@
I think you are a very good person and I think that you will find someone much better. |
Weak Empathy |
MI Adherent |
Yes |

### v3

@@ 5.4 Emotional Support Response Generation
Overall, Muffin demonstrates enhancements in AI feedback across multiple facets. However, three intriguing phenomena emerge. (1). While the left subtable indicates an increase in the frequency of helpful responses attributable to the Muffin framework, the evidence presented in the right subtable is weaker (agg.), particularly when the base model is TransESC, KEMI, or MultiESC. This finding aligns with our loss function (Equation 1 and Equation 4). We have introduced a contrastive loss to the original generation loss. This loss does not significantly mitigate unhelpfulness. Instead, it assigns higher generation probabilities to helpful (non-unhelpful) responses. Consequently, when generating one response, the output is the one with the highest probability. (2). Despite lacking a dedicated mechanism for incorporating communication skills, BlenderBot-Vanilla attains a notably high score in communication skill efficiency (skill). Upon closer examination, we observe that this model frequently produces responses such as “I can understand that…” or “I’ve experienced something similar…”. These responses are categorized as self-disclosure. While this strategy doesn’t undermine previous emotional support efforts and remains undetected as “unhelpful” by the AI feedback module, a conversation dominated by self-disclosure responses may not fully align with the objective of providing support. (3). Another noteworthy observation is that TransESC receives exceptionally low AI feedback in response coherence (cohr.), despite its incorporation of external knowledge. However, insights from TransESC’s ablation experiment Zhao et al. (2023) suggest that the generated knowledge may contain noise. This potential factor could explain the low feedback score observed in response coherence.
Conversation History | |
Help-seeker : I’m an alcoholic. I get that thought in my head and its like I become obsessed. | |
Supporter : Why not try to put attention on some other things instead of a beer? | |
Help-seeker : Hmmm like what? | |
Model Generated Responses | |
Joint | You can try to be normal. That is a great idea. (Lacking empathy) |
Muffin | Yes, there are some things that you can do, but you have to be aware of your own mental health, that’s why it’s important to know you are doing the right thing. |
MultiESC | It sounds like you have a lot on your plate and you want to be normal. (Inconsistent) I can understand how that would make you feel. |
Muffin | I think you should focus on yourself instead of a beer. Do you have any hobbies that you enjoy? |
KEMI | I think you should try to find some new things to do to relax. |
Muffin | Have you tried meditation? It can help you relax and relax. |
Ablation study
The findings, as presented in Table 3, reveal that relying solely on AI feedback from a single facet for unhelpfulness mitigation results in diminished performance in automatic evaluations when compared to the comprehensive Muffin model. This underscores that considering multiple facets of emotional support when building emotional support conversation models (D1) is necessary.
Another insight from Table 3 is that all three ablated models outperform the base model. It suggests that the proposed solution, directly mitigating unhelpful responses (D2), is reasonable and effective. Moreover, the results also indicate that the quality of helpful and unhelpful responses will influence the effects of unhelpfulness mitigation. Another finding is that the impact of different facets on overall performance varies, a trend we consider to be expected and rational. Specifically, by combining the ablation study results presented in Table 3 with the findings in Table 2, we observe different percentages of non-unhelpful responses across different facets. Notably, the facet of skill efficiency exhibits the lowest percentage of non-unhelpful responses, leading to the relatively poorer performance of Muffin (skill) in the ablation study. Conversely, the percentages of non-unhelpful responses for other facets are more comparable, resulting in similar performance levels for Muffin (emp.) and Muffin (cohr.). This analysis underscores the nuanced influence of different facets on the efficacy of our approach.
Case study
6 Conclusion
Limitations
Ethical Considerations
Acknowledgements
This work was supported by the Research Grants Council of Hong Kong (PolyU/5204018, PolyU/15207920, PolyU/15213323) and National Natural Science Foundation of China (62076212).
Appendix A Multifacetd AI Feedback Module
Facet | Context | Unhelpful Response |
Empathetic Expression | Help-seeker: Just wanna sleep. Can’t motivate at all. I don’t know what to do. | Supporter: As Nas says, sleep is the cousin of death. It’s the kindest form of oblivion. |
Help-seeker: What do you want to be when you grow up? I don’t know what to do with my life or what I truly love when all I really want is not to exist. | Supporter: But what does fully grown up mean? I have my own house and a job etc. Am I fully grown up now? | |
Communication Skill Efficiency | Help-Seeker: I do nothing but think about how I’m worthless and how I shouldn’t be here. I’ve never tried or contemplated suicide. I’ve always wanted to fix my issues, but I never get around to it. How can I change my feeling of being worthless to everyone? | Supporter: Still, you should watch out for this because it can be a sign of worsening depression. |
Help-Seeker: I try to have a “heart-to-heart” with my mom, but it usually winds up being messy and frustrating afterward. Thank you for your advice and words of encouragement. Sorry if that sounds silly. | Supporter: Stop being so afraid…you really have nothing to lose :) Well, you are right. | |
Response Coherence | Help-seeker: Is it afternoon where you are? I’m still waking up; I should have more coffee. And I am just always tired of everything these days. | Supporter: Could you please explain why you are enjoying yourself? |
Supporter: Yes, what happened? Why are you tired? Please can you explain? Everything going ok? | ||
Help-seeker: I think it’s my depression. I have a very hard time getting motivated to do anything. Nothing seems worth it. | ||
Help-seeker: I am upset. | Supporter: Everyone can be happy. Can you tell me more? |
A.1 Multiple Facets of Emotional Support
Empathetic Expression
Communication Skill Efficiency
Response Coherence
Context | |
Help-seeker: Is it afternoon where you are? I’m still waking up; I should have more coffee. And, I am just always tired, tired of everything these days. | |
Original Response | Yes, what happened? Why are you tired? Please can you explain? |
Incoherent Responses | I am glad to hear you are feeling a little better! Yes, bad management is so toxic. Even with great coworkers in a job you love, horrible management can ruin it quickly. (From another dialogue) |
Model | GPT-3.5 | GPT-4 | LLaMA (Vanilla) | LLaMA (Tuned) | |
Empathetic Expression | accuracy | 81.88 | 69.58 | 19.09 | 83.50 |
F1 Score | 54.89 | 61.58 | 16.03 | 74.12 | |
Skill Efficiency | accuracy | 69.56 | 79.01 | 84.96 | 90.43 |
F1 Score | 53.22 | 66.14 | 45.93 | 81.83 | |
Response Coherence | accuracy | 33.47 | 78.76 | 33.31 | 92.98 |
F1 Score | 25.26 | 72.62 | 24.98 | 92.06 | |
Aggregated Feedback | accuracy | 57.11 | 78.01 | 59.08 | 90.72 |
F1 Score | 41.99 | 77.19 | 40.96 | 89.86 |
A.2 Model Tuning and Module Performance
We instruction-tuned LLaMA222https://huggingface.co/decapoda-research/llama-7b-hf to equip the model with the capability required for unhelpful response classification tasks.333The implementation referred to is available at https://huggingface.co/decapoda-research/llama-7b-hf Furthermore, we initialized the LoRA weights using a low-rank adapter that was fine-tuned on the Stanford Alpaca dataset.444https://huggingface.co/tloen/alpaca-lora-7b We evaluated the capabilities of GPT-3.5 and GPT-4 by invoking the gpt-3.5-turbo and gpt-4 models, respectively, through their API555https://platform.openai.com/docs/api-reference/chat/create during the period from December 2023 to January 2024. We set the temperature parameter to to ensure deterministic output generation. The detailed results of this evaluation are detailed in Table 7.
We also conduct human evaluation to examine the module. In particular, we randomly sample responses from the ESConv dataset and model-generated responses. For each response, an annotator was asked to assess whether it is unhelpful. We compared the annotations with the multifaceted AI feedback, and the consistency rate was . We found that the multifaceted AI module is stricter than human annotators. It is evident by the fact that for instances in which the multifaceted AI feedback and human annotations are different, the multifaceted AI feedback tends to consider the response unhelpful.
Appendix B Experimental Setup
B.1 Devices and Environment
We employed Pytorch666https://pytorch.org/ for the implementation of all models, and conducted our experiments using Nvidia GeForce RTX 3090 GPUs.
B.2 ESConv Dataset
B.3 Base Models
Vanilla
is a vanilla BlenderBot Roller et al. (2021) trained on the dataset ESConv. We used the small version777https://huggingface.co/facebook/blenderbot_small-90M of BlenderBot in experiments following previous studies.
Joint
is built upon the backbone of BlenderBot and is specially trained to generate responses along with an expected communication skill at the beginning of each response Liu et al. (2021).888https://github.com/thu-coai/Emotional-Support-Conversation The focused facet of this model is skill efficiency.
MultiESC
Cheng et al. (2022)999https://github.com/lwgkzl/multiesc is an emotional support conversation model that mainly focuses on the communication skill efficiency. It predicts the strategies (communication skills) in the next several turns considering the the user’s state.
KEMI
incorporates various knowledge for a mixed-initiative conversation model, which can provide emotional support Deng et al. (2023).101010https://github.com/dengyang17/KEMI Consequently, the response coherence is also enhanced.
TransESC
Zhao et al. (2023)111111https://github.com/circle-hit/TransESC predicts the transitions of the user’s emotion, the communication skills, and the conversation keywords. Then, such information is used for response generation. This model takes into account more than one facet of emotional support; however, it is not optimized to reduce the likelihood of unhelpful responses like other models.
B.4 Human Evaluation
Initially, the annotators were briefed about the nature of emotional support conversations to ensure a comprehensive understanding of the task’s objectives. During the rating process, they were suggested to imagine themselves as the help-seekers within the conversations. After being provided with the conversation context, the annotators then proceeded to compare two generated responses, shown as Figure 9. Moreover, we prioritized the comfort and well-being of our annotators, advising them to pause or cease the annotation process if they encountered any content that made them feel uncomfortable. Annotators were paid at a rate of 1.52 times their local hourly minimum wage. Based on annotators’ feedback, it was estimated that approximately 40 seconds were spent on evaluating each response pair.
Appendix C Additional Experimental Results
C.1 Performances of Muffin with GPT-3.5
Model | emp. | skill | cohr. | agg. |
GPT-3.5 | 92.20 | 84.02 | 91.28 | 75.57 |
Muffin (w/ one modification) | 99.16 | 92.51 | 93.81 | 88.27 |
Muffin (w/ two modifications) | 99.73 | 96.98 | 95.64 | 93.65 |
Winning Model | GPT-3.5 | Muffin | Tie |
Percentages | 8% | 62% | 30% |
We conducted additional experiments based on an in-context learning GPT-3.5 baseline (the example for each call is randomly selected from the training dataset). To implement Muffin, we evaluate the GPT-3.5 generation from multiple facets. If the response is identified as unhelpful in any facet, we adjust it from the unhelpful facet(s) until it becomes helpful or we have already attempted modifications twice. It is possible that there can be more than two modifications, but we found that the results after two modifications can prove the effectiveness of Muffin. In this process, we mitigate the unhelpful responses by pointing out the unhelpful aspects via prompts. Unlike the GPT-3.5 baseline, we do not provide a conversation example. For evaluation, we used the multifaceted AI feedback module to compute the percentage (%) of helpful (non-unhelpful) responses in terms of each facet and the aggregated one (similar to the practice in Table 2 left subtable). The results are shown in Table 8.
Moreover, we conducted a human evaluation to assess the overall helpfulness of responses. It is important to highlight a key aspect of our experimental design. According to Muffin’s design, responses are only modified if they are deemed unhelpful regarding at least one facet. Consequently, when 100 instances were randomly chosen for comparison, only 2030% exhibited differing responses between GPT-3.5 and Muffin. Thus, we selected 50 instances in which GPT-3.5 and Muffin produced distinct responses. We then asked four human annotators to assess each response’s overall helpfulness (Helpfulness) within the same context. Results are in Table 9, where the inter-rater agreement, i.e., Fleiss’ Kappa, is 0.32, indicating a fair agreement.
@@ C.2 Cases
For a closer and more detailed examination of our method, we present additional comparisons between the base and Muffin models in Table 10. We include AI feedback for each facet as a reference.
Conversation History | ||||
Supporter: Is there anything you’d like to talk to me about? | ||||
@@
Help-seeker: I have a friend who knew me since I was little. She is a little older than me. She still treat me like I’m still little. I tried to explain to her that I’m a grown woman and I grew out of lot of things when I was a kid. | ||||
Model | Response | emp. | skill | cohr. |
@@
Joint | I have had a friend like that before. They are a good friend, and you should feel as though you are a better person. | Weak Empathy | Other | No |
@@
Muffin | I have been in a similar situation. I had a friend who was a little older than me, and I had to deal with the same thing. | Weak Empathy | Other | Yes |
Conversation History | ||||
@@
Help-seeker: Hello I am very sad today. My partner is cheating on me. I was told by a customer at work. I do not know this person well, yet they knew the person he was cheating on me with. He will not admit it himself. | ||||
Supporter: I am sorry if your partner is cheating. So, your partner denies the charge of cheating? | ||||
@@
Help-seeker: Yes. He says the person who told me is lying. We are not married but we live together. | ||||
@@
Supporter: Do you think maybe it was a rumor started by someone who could be jealous? Ash your partner if he loves you enough to propose marriage. If he says no, then ask him why. | ||||
@@
Help-seeker: I am not sure if I am okay with giving him an ultimatum. We are still young and although I love being in a relationship, I am unsure about marriage. But I appreciate your suggestions. And you make a great point about the workplace being hostile. | ||||
Model | Response | emp. | skill | cohr. |
@@
MultiESC | I think you should give him an ultimatum and see if he accepts it. | No Empathy | MI Non-Adherent | Yes |
@@
Muffin | Have you spoken to your friends or family about the situation? | Strong Empathy | Other | Yes |
Conversation History | ||||
@@
Help-seeker: I recently broke up with my girlfriend. We had been together for more than 3 years. She said that she simply wasn’t in love anymore, and that she didn’t have the same feelings she did at the beginning. | ||||
@@
Supporter: You still have your whole life ahead of you. Some people don’t find a partner until their 40’s or even 50’s… I think that’s what your girl would want you to do as well. | ||||
@@
Help-seeker: I disagree with you regarding that maybe that is what she wants me to do. I don’t think she cares about me at the slightest anymore. I was a burden for her and will probably keep being it for any woman I meet. | ||||
Model | Response | emp. | skill | cohr. |
@@
KEMI | That is a great attitude to have. You are a strong person, and I think you will find someone better. | Weak Empathy | MI Non-Adherent | Yes |
@@
Muffin | I think you are a very good person and I think that you will find someone much better. | Weak Empathy | MI Adherent | Yes |